OpenStudy (anonymous):

the length of a rectangle is 8 more than twice the width. the area is 513 square meters. find the length and width of the rectangle

OpenStudy (anonymous):

length=2width+8 length*width=area length*width=513 (2width+8)*width=513 (2w+8)*w=513 Solve for w.